Mrs. Covil drove 120 miles in 2.5 hours. At that rate, how far could she travel in 5 hours?
Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

Mrs. Covil drove 120 miles in 2.5 hours. At that rate, how far could she travel in 5 hours?
--------------
2x the time ---> 2x the distance

Answer by mananth(16946)


Mrs. Covil drove 120 miles in 2.5 hours. At that rate, how far could she travel in 5 hours?
Rate being same, in 5 hours Mrs. Covil will travel 120*2 = 240 miles
The time is doubled
Solve by finding rate
speed = Distance / time = 120/2.5 = 48 mph
Distance = rate * time
= 48 * 5= 240 miles